OpenStudy (vjt):

Identify the correct x intercept and y intercept of the equation x - 2y = 8. x intercept = (8, 0); y intercept = (0, 4) x intercept = (8, 0); y intercept = (0, -4) x intercept = (-8, 0); y intercept = (0, -4) x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (gorv):

y intercept means x=0 put in equation 0-2y=8 -2y=8 divide both side by -2 y=8/-2 y=-4 this is y intercept

OpenStudy (gorv):

x intercept means y=0

OpenStudy (gorv):

x-0=8 x=8 means x intercept

OpenStudy (vjt):

thanks